# Pay Attention: a Call to Regulate the Attention Market and Prevent Algorithmic Emotional Governance

### v1

@@ 7 Thank you for… your attention
AI is domain-independent. It is being applied in all our areas of interest: information, business, money, politics, employment, sports, games, sex etc. And the worldwide deployment of these techniques, partly due to its coupling with the Web, could have detrimental consequences in all these areas alike, unless properly regulated. This is a commonplace observation but it is the reason why, to prevent such detrimental effects, an ethical AI approach to AI governance must be multidisciplinary and interdisciplinary.
With this mindset, this paper brought together conclusions from more than 70 articles and books from different disciplines (psychology, sociology, neuroscience, politics, legal domain, computer science, education, etc.) to analyze and call for actions against the current practices competing to capture our attention in several “Web Wild West corners”. The problem is both critical and complex, and authors of [34] defend the need for a “nuanced multidimensional view of how social media use may shape information consumption” and they urge us to consider “the complex variety in social media platforms [and the] considerable variation in observed impacts among them”. In  [22] authors add that “This research requires to navigate the careful tension between privacy, security, economic interests, censorship and cultural differences, and requires to be addressed from multiple disciplines that can assess not only the technological aspect, but also the individual and the social one (…) There is ample room for investigation (…), opening a novel, exciting and interdisciplinary line of research.”
@@
Recommendation systems that learn to predict us effectively learn to manipulate us, and to be predictable is to lose freedom. Everyday we fuel the predictors in exchange for immediate satisfaction and instant pleasure, this amounts to continually mortgaging our freedom. Besides, these systems that compete for our attention end up pressuring us to consume and to react more and more quickly to their recommendations. And, as we know, acceleration is a form of alienation [53].
In another context and to address our own human limitations, [10] recommended to find ways to increase our overall level of consciousness and reclaim the power of long-term reflection. Our leaders and role models [^15] A role model is a person whom others look at as an example to be imitated. struggle to embody the values of patience, conscience and moderation [10], but our computer systems rarely drive us in that direction either. On the contrary, current AI applications are pushing us not to use our conscience, but to play their automation game. Yet there is no reason for these systems to live in our mind rent free and it is urgent to redesign them so they regularly push us to take a step back, to be more conscious of what we are doing, viewing, saying, spreading, etc. The challenge is to (re)take and (re)give time for awareness, attention and reflection: we need to (re)take that source of freedom. And for this, we proposed a non-exhaustive first set of principles to (re)design Web applications and inscribe in them a set of agreed-upon values.

### v2

@@ 7 Thank you for… your Attention
AI is domain-independent. It is being applied in all of our daily life areas of interest: information, business, money, politics, employment, sports, games, sex etc. And the worldwide deployment of these techniques, partly due to its coupling with the Web, could have detrimental consequences in all these areas alike, unless properly regulated. This is a commonplace observation but it is the reason why, to prevent such detrimental effects, an ethical AI approach to AI governance must be multidisciplinary and interdisciplinary.
With this mindset, this paper brought together conclusions from more than 70 articles and books from different disciplines (psychology, sociology, neuroscience, politics, legal domain, computer science, education, etc.) to analyze and call for actions against the current practices competing to capture our attention in several “Web Wild West corners”. The problem is both critical and complex, and authors of (Kitchens, Johnson, and Gray 2020) defend the need for a “nuanced multidimensional view of how social media use may shape information consumption” and they urge us to consider “the complex variety in social media platforms [and the] considerable variation in observed impacts among them”. In  (Fernández, Bellogín, and Cantador 2021) authors add that “This research requires to navigate the careful tension between privacy, security, economic interests, censorship and cultural differences, and requires to be addressed from multiple disciplines that can assess not only the technological aspect, but also the individual and the social one (…) There is ample room for investigation (…), opening a novel, exciting and interdisciplinary line of research.”
@@
Recommendation systems that learn to predict us effectively learn to manipulate us, and to be predictable is to lose freedom. Everyday we fuel the predictors in exchange for immediate satisfaction and instant pleasure, this amounts to continually mortgaging our freedom. Besides, these systems that compete for our attention end up pressuring us to consume and to react more and more quickly to their recommendations. And, as we know, acceleration is a form of alienation (Rosa and Chaumont 2017).
In another context and to address our own human limitations, (Bronner 2021) recommended to find ways to increase our overall level of consciousness and reclaim the power of long-term reflection. Our leaders and role models [^13] A role model is a person whom others look at as an example to be imitated. struggle to embody the values of patience, conscience and moderation (Bronner 2021), but our computer systems rarely drive us in that direction. On the contrary, current AI applications are pushing us not to use our conscience, but to play their automation game. Yet there is no reason for these systems to live in our mind rent free and it is urgent to redesign them so they regularly push us to take a step back, to be more conscious of what we are doing, viewing, saying, spreading, etc. The challenge is to (re)take and (re)give time for awareness, attention and reflection: we need to (re)claim that source of freedom. And for this, we proposed a non-exhaustive first set of principles to (re)design applications and inscribe in them a set of agreed-upon values.